You buy 6 bags of gummy bears and 3 bags of chocolate for a total of $24.75. Your friend decides to buy 2 bags of gummy bears and 5 bags of chocolate for a total of $21.25.  How much do the bags of chocolates cost?
Answer by rothauserc(4718)

let g be the cost of a bag of gummy bears and c be the cost of a baag of chocolate
:
1) 6g +3c = 24.75
:
2) 2g +5c = 21.25
:
solve equation 1 for c
:
3c = 24.75 -6g
:
c = 8.25 -2g
:
substitute for c in equation 2 and solve for g
:
2g +5(8.25 -2g) = 21.25
:
2g +41.25 -10g = 21.25
:
-8g = -20.00
:
g = 2.50
:
c = 8.25 -2(2.50) = 3.25
:
****************************************************
a bag of chocolate costs $3.25
:
check answer
:
equation 1
:
6(2.50) +3(3.25) = 24.75
:
15.00 +9.75 = 24.75
:
24.75 = 24.75
:
equation 2
:
2(2.50) +5(3.25) = 21.25
:
5.00 +16.25 = 21.25
:
21.25 = 21.25
:
answer checks
********************************************************
: